OpenStudy (darkprince14):

What is the probability that a person flipping a fair coin requires four tosses to get a head?

OpenStudy (perl):

This is the probability of getting tails on first three tosses, then heads.

OpenStudy (perl):

P(TTTH)

OpenStudy (darkprince14):

Therefore, the answer is 1/16. But the answer on the book is 1/4. I don't quite get that.

OpenStudy (darkprince14):

A similar question was also with it. The problem where you will have exctly one heads after 4 tosses. The answer indicated in the book was 1/16 which is P(TTH)

OpenStudy (perl):

I'm not sure why it does not work.

geerky42 (geerky42):

Are you sure? Because these answer should be the other wat\y.

geerky42 (geerky42):

way*

OpenStudy (darkprince14):

the second question actually wasn't exactly like that.. Here is the second question whose answer indicated in the book was 1/16 "Suppose an experiment consists of tossing a fair coin until three heads occur. What is the probability that the experiment ends after exactly six flips of the coin with a head on the fifth toss as well as on the sixth" The answer for the 1st question was 1/4 (indicated). It only showed the answer, no solutions.

geerky42 (geerky42):

You have trouble understanding why answer for second question is 1/16? And what is first question exactly?

OpenStudy (darkprince14):

the probability of getting a head after four tosses... was just comparing the two of them since i still don't know how to differentiate the two.

OpenStudy (darkprince14):

**the probability that the first head will occur on the 4th toss of the coin

OpenStudy (amistre64):

P(H|TTT) = P(H n TTT)/P(TTT) seems kinda pointless eh

OpenStudy (amistre64):

it could be that your book is just full of errors :)

OpenStudy (darkprince14):

I used the pdf for geometric distribution in the (1st) question and got 1/16 as well. Was wondering why I got the wrong answer.

OpenStudy (rational):

first question is geometric distribution. first head at fourth toss means the first three tosses are tails : 1/2*1/2*1/2 and the last toss is head : 1/2 since each trial is independent, probability of getting a head at fourth toss would be 1/2^4 = 1/16 textbook answer is wrong

OpenStudy (darkprince14):

so I should just stick to 1/16 then? I have an exam later and this helped a lot :) thanks for the enlightenment and for the concepts you guys taught me :D

OpenStudy (rational):

Yes 1/16 is right

OpenStudy (rational):

`Suppose an experiment consists of tossing a fair coin until three heads occur. What is the probability that the experiment ends after exactly six flips of the coin with a head on the fifth toss as well as on the sixth` first fix fifth and 6th results. probability of occuring heads on fifth and sixth tosses is 1/2*1/2 = 1/4